Solve.
6(a + 3) = 18 + 6a
(Points : 1)
no solution
all real numbers
–3
–1

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Solve.
6(a + 3) = 18 + 6a
6a+18 = 18 + 6a
--------
Solution:: True for all Real Numbers.
------------
Cheers,
Stan H.
---------------

Answer by ewatrrr(24785)

6(a + 3) = 18 + 6a
6a + 18 = 18 + 6a
0 = 0
All real Numbers the Solution